My teacher is not very good at explaining the lessons to me.
I do not get the concept of dividing exponents.
PLEASE HELP ME!

Answer by stanbon(57953)

You do not divide exponents; you add or subtract them.
For example 3^5/3^2 = 3^(5-2) = 3^3
OR
3^5 * 3^2 = 3^(5+2) = 3^7
-----------
Hope this helps.
Cheers,
Stan H.